OpenStudy (anonymous):

The Doe family is ready to fill their new swimming pool. It can be filled in 12 hours if they use their own water hose, and in 30 hours if they use Mr. Jones', their neighbor's water hose. How long will the Doe's take to fill their pool if the neighbor's hose is used along with their own?

OpenStudy (retireed):

I believe I have solved similar problems by adding the inverses of the time for example... 1/12 + 1/30 = 1/ x x will equal both hoses used together Can you solve this equation for X ?

OpenStudy (retireed):

You can prove this works because say you have two hoses of equal flow rate.... one hose 10 hours and another equal hose at 10 hours 1/10 + 1/10 = 2/10 or 1/5 both combined will take 5 hours